# Building a Scalable, Effective, and Steerable Search and Ranking Platform

### v1

@@ 1. INTRODUCTION
The ability to provide a personalized, dynamic, scalable, and efficient ranking platform that can be employed across various experiences has become critical to driving customer engagement and business value in e-commerce. We achieve this by building a ranking platform grounded in four key design principles: 1) composability - our platform consists of multiple state-of-the-art ranking models and candidate generators working in orchestration; 2) scalability - ensured by vector-based indexing and scoring; 3) shared real-time serving infrastructure and 4) steerable ranking that can adapt to varying customer preferences and business objectives.
@@
Our platform is able to support the integration of multiple models through vertical layering, and horizontal integration, blending the outputs of various models or other candidate sources. This enables scalability, independence, and ability to mix various content types to cater to specific use cases as well as building ranking ensembles by combining the outputs of multiple models. The platform’s scalability is driven by employing a vector store in its candidate generation stage, facilitating efficient indexing, scoring, and retrieval which are crucial for managing a growing catalog of items and number of customers. To this end, we compute dense representations of customer behaviors, contexts, and item inputs in a common embedding space. Built on a foundation that allows for near real-time scoring and computation of customer and item representations, our platform dynamically adapts the ranking in all layers to recent customer changes (including candidate generation) which increases the probability of discovering relevant items (Xia et al., 2023; Qi et al., 2020; Chen et al., 2019). Moreover, utilizing efficient transformer-based model architectures across all layers allows sharing of the serving infrastructure, which in turn reduces engineering complexity, increases re-usability, and helps avoid inconsistency between training and serving phases, which is a common issue in machine learning engineering systems (Zinkevich, 2010).
@@
Many ranking systems in the literature rely on pre-trained items and customer embeddings. Our experiments reveal that similarly to NLP tasks (Devlin et al., 2019), the effectiveness of our models significantly increases when the pre-trained input item embeddings are further fine-tuned on the ranking task. Notably, we show that if these embeddings are not continuously trained, the candidate generation model shows substantially less customer engagement. To address the item cold-start problem, we introduce epsilon-greedy exploration by blending fresh items from additional candidate sources into the organic ranking. We address the customer cold-start by leveraging customer context and in-session capabilities and data.
@@ 2. RELATED WORK
Thanks to their advantages over traditional deep-learning-based models, sequence-based recommender systems (Kang and McAuley, 2018; Sun et al., 2019) in their two flavors of language modeling (CLM and MLM) (de Souza Pereira Moreira et al., 2021; Petrov and Macdonald, 2023) have gained wide traction. These systems have proven powerful in modeling customer behavior as a sequence of actions due to their capability to 1) capture both short-term and long-term interests (Xia et al., 2023); and 2) the ability to compute complex feature interactions. However, most existing works use public datasets that in certain cases are not even adequate for sequential recommendation tasks (Hidasi and Czapp, 2023; Kang and McAuley, 2018; Sun et al., 2019; de Souza Pereira Moreira et al., 2021) and many of focus only on offline experiments, with only a few works reporting actual customer impact through end-to-end A/B testing in large scale environments (Xia et al., 2023; Pancha et al., 2022). Our paper extends this line of work and demonstrates the usefulness of these models in real-world applications that include personalized item browsing and search. Moreover, to the best of our knowledge, this is the first work to apply transformer networks as part of the two-tower architecture for learning embeddings for ranking use cases that were originally introduced in the line of work of Google (Covington et al., 2016; Yang et al., 2020). Unlike related approaches (Chen et al., 2019; Xia et al., 2023), our approach to scoring candidate items does not suffer from performance issues caused by a long sequence length (Chen et al., 2019) or limiting the input embedding size due to concatenating the candidate item embeddings with the action sequence embeddings. Unlike (Xia et al., 2023), we do not observe a significant drop in ranking performance and feed only the average embedding of the candidates as a fixed input position into the transformer network.
Moreover, only a few published studies (Steck et al., 2021; Celikik et al., 2022a; Chen et al., 2019; Xia et al., 2023; Celikik et al., 2022b) include heterogeneous inputs such as context and content-based features, which help address the cold-start problem and address data sparsity by improving generalization. For example, (Chen et al., 2019) considers customer profiles and contextual features by using wide & deep learning (WDL), which relies on a concatenation of signals in the output of the network, making it inadequate to capture powerful feature interactions. (Xia et al., 2023) employs deep and cross-network (DCN) on top to explicitly model feature-crosses, which significantly increases the number of parameters of the network.
@@ 3. SYSTEM DESIGN
Designing a highly performant, scalable, and steerable ranking platform entails multiple challenges and complex choices. The existing literature often focuses only on subsets of them (Lakshmanan et al., 2020; Covington et al., 2016; Yang et al., 2020; Essinger et al., 2021), while this section aims to navigate through them holistically. We describe foundational design principles and provide an overview of our system architecture and components (see Figure 2). The presented design is generalizable and applicable to other retrieval and ranking use cases and setups.
@@
Composability and orchestration of multiple models. Our platform supports the integration of various models either ”vertically” for layered ranking and retrieval, or ”horizontally” by blending outputs from different models or candidate generators. It features a multi-layered architecture that enhances scalability and reduces inter-dependencies. The initial layer retrieves relevant candidates from multiple candidate generators, possibly generating different item or content types for use cases such as feeds. Each candidate generator typically entails a (lightweight) ranking model. Subsequent layers refine these selections; the ranking layer applies heavy personalized models for ranking pre-selected items of possibly different types. The policy layer ensures compliance with business or product specifications. Blending strategies mix outputs to suit specific needs, such as combining different content types in desired proportions (e.g. product, outfits, videos) or balancing popular, fresh, and personalized content (Gomez-Uribe and Hunt, 2016). Model blending is also the blueprint for building ranking ensembles, where outputs of multiple models are combined either by score weighting or meta rankers.
@@
Scalable platform. The multi-layered model architecture also allows us to achieve high scalability. The more accurate, but computationally heavier, part of the ranking is performed on later layers only on a small subset of candidates obtained from the less accurate but more computationally efficient candidate generator layers. Thus, the highest scalability requirements are placed on the candidate generator. Scalability of the candidate generator layer is achieved through a vector store allowing efficient indexing, scoring, and personalized retrieval capabilities to scale for a large and growing item catalog and customer base. Our platform includes infrastructure to compute customer vectors based on past actions and context (such as country, browsing category or search query, etc.) and item vectors. The item vectors are indexed in an internally hosted vector store that supports efficient approximate nearest neighbors search and retrieval.
Shared near real-time serving infrastructure. Similarities in model types across layers allow for shared training datasets and much of the serving infrastructure, enhancing efficiency. Deep learning minimizes feature engineering complexity by integrating embedding mappings directly into the model graph, ensuring consistency between training and serving phases. Shared online feature stores across ranking layers enable effective input caching for each request, simplifying engineering efforts.
@@
Steerable ranking. Our system’s flexibility allows for external adjustments to ranking objectives to align with business goals like customer satisfaction or profitability. It also supports diverse content types through its candidate generators and mixing components and integrates business heuristics in the policy layer.
In the following sections, we present in more detail the main components of our multi-layered system, i.e., the candidate generator layer, the ranking layer, and the policy layer.
4. CANDIDATE GENERATION
@@
The objective of our candidate generation layer is to generate personalized item candidates from the item catalog for each individual customer efficiently in near real-time. According to our findings, a personalized candidate generator is essential for the performance of the overall ranking system (see section 8 for details).
@@
We follow the classical two-tower approach (Yi et al., 2019), where the customer tower processes historical customer action sequences and contextual data to generate a customer embedding while the item tower is responsible for generating item embeddings. These embeddings are then combined by using dot product to generate a score per item as shown in Figure 3 (theoretical justification about the expressiveness of the two-tower model is provided in the Appendix, subsection 10.4).
Although trained together, the towers are deployed and operated independently. The item tower generates item embeddings that are indexed in a vector store for efficient ranking and retrieval. The customer tower is invoked to generate customer embeddings each time a customer accesses the platform. The freshly computed customer embeddings are then used to find items with similar embeddings in the nearest-neighbors index.
@@
To compute a customer embedding from the customer action sequence and the context, we employ a transformer encoder and use causal language modeling (CLM) as in (Kang and McAuley, 2018), processing each customer sequence once per epoch. We predict the subsequent item in the sequence while preventing backward attention by using a causal mask.
For simplicity, items in the item tower in Figure 3 are represented by using a single embedding that jointly encodes product metadata (brand, category, material, etc.) and visual cues. It is worth mentioning that in this common setting, trainable input embeddings on the ranking task performed substantially better in both offline and online tests. Details on how different types of input signals from the customer journey data (e.g. contextual and customer action sequence data) are encoded in the customer tower of the model are captured in Section 5.2.
@@ 5. RANKING LAYER
The objective of the ranking layer is to rank items returned by the candidate generation phase by their relevance to the customer by using a powerful ranking model. We model this task as a pointwise multi-task prediction problem, where we predict the probability of the customer performing any of the following positive actions on a candidate item: click, add-to-wishlist, add-to-cart, purchase given the context and their past behavioral data. If a candidate item is associated with any of these positive actions, we consider it a positive item, otherwise, it is considered a negative.
@@ 5.1. Model Architecture
Figure 4 depicts the architecture of the model in the ranking layer. It consists of four main parts: embedding layer, item candidate embedding, customer and context embedding computed via a self-attention mechanism, a prediction head for each of the target action types, and a shallow position branch used for position debiasing.
@@
For each of the target actions, we define a prediction head, which takes customer and candidate item representations as inputs. The score, for a given target action, is obtained by computing a dot product between the customer and context embedding and all candidate item embeddings in parallel, after passing them through a FFN. A sigmoid function is used to normalize the score and interpret it as a probability. During training, each prediction head contributes equally to the loss, while at serving we produce the final ranking by weighting the scores of each prediction head. The weights are dynamically configurable and determined analytically depending on the customer touch point.
@@
The training objective of our main candidate generator is based on predicting the next customer action, however, our data pipelines are configurable and can be easily extended to predict actions in a longer future time window, e.g. a week, to balance long and short-term customer preferences and improve diversity in outputs.
@@ 5.2. Encoding of the Customer Journey Data
In this section, we elaborate on how we encode holistic customer journey data into models in both layers. This consists of (1) behavioral data which includes customer action sequences, item and action metadata, and temporal data, and global and local contextual data. We argue and demonstrate in our experiments that providing complete and heterogeneous information to the model is crucial to predictions that are personalized and contextually relevant.
Behavioural data. As already mentioned, customer action sequences are encoded using transformer encoders as these have been proven more effective than other approaches (Xia et al., 2023). Each action in an action sequence is represented by item embedding that encodes domain-specific visual information, categorical item metadata, action type, and timestamp embeddings. The categorical item metadata consists of relevant attributes such as brand, color, pattern, category, material, etc. The timestamp embedding encodes quantized timestamps as measured from the beginning of the model training. Temporal data is crucial for modeling customer behavior across sessions. Since customer intent can vary drastically across different sessions, modeling action sequences while ignoring this structure affects performance negatively (Rahmani et al., 2023; Celikik et al., 2022b). All inputs are passed through trainable embedding to project discrete or bucketed values into low-dimensional spaces. Item and action-specific embeddings are concatenated.
@@ 5.3. Position Debiasing
In the context of ranking systems, feedback loops occur when a model influences customer interactions which can lead to biased relevance data. Typically, items ranked higher by the model receive more customer attention, causing position bias. Position bias causes a skewed representation of actual customer preferences which in turn may amplify and degrade model performance over time due to the feedback loop. To address this, we incorporate position information as a feature into a model to separate the effect of the position and the true relevance of the probability with which the customer would interact with an item. A debiased model is conditioned on positions during training (right branch in section 5) and position-independent during serving by setting positional feature to fixed values to counter position bias (Zhao et al., 2019). The position branch is separate from the rest of the model due to the asymmetry between training and serving. We performed additional A/B test which confirmed that adding position debiasing resulted in increasing long-tail utilisation by 5% while not deteriorating engagement metrics.
@@ 6. POLICY LAYER
The last stage of the system, the policy layer, is responsible for the final page composition. Here, multiple re-ranked candidate items are combined into one, performing granular, page-level optimization, and applying heuristics, business rules, and filters depending on the use case. In the following, we describe how we promote fresh (new or cold-start) items while simultaneously introducing exploration into the system. We also describe some common heuristics applied in this layer to meet product requirements.

### v2

@@ 1. INTRODUCTION
The ability to provide a personalized, real-time, and scalable ranking platform that can be employed across various experiences has become critical to driving customer engagement and business value in e-commerce. We achieve this by building a ranking platform grounded in four key design principles: 1) composability - our platform consists of multiple state-of-the-art ranking models and candidate generators working in orchestration; 2) scalability - ensured by vector-based indexing and scoring; 3) shared real-time serving infrastructure and 4) steerable ranking that can adapt to varying customer preferences and business objectives.
@@
Our platform is able to support the integration of multiple models through vertical layering, and horizontal integration, blending the outputs of various models or other candidate sources. This enables scalability, independence, and ability to mix various content types to cater to specific use cases as well as building ranking ensembles by combining the outputs of multiple models. The scalability of the first or candidate generation layer is ensured by employing a vector store, facilitating efficient indexing, scoring, and retrieval which are crucial for managing a growing catalog of items and number of customers. To this end, we compute dense representations of customer behaviors, contexts, and item inputs in a common embedding space. Built on a foundation that allows for near real-time scoring and computation of customer and item representations, our platform dynamically adapts the ranking in all layers to recent customer changes which increases the probability of discovering relevant items (Xia et al., 2023; Qi et al., 2020; Chen et al., 2019). Moreover, utilizing efficient transformer-based model architectures across all layers allows sharing of the serving infrastructure, which in turn reduces engineering complexity, increases re-usability, and helps avoid inconsistency between training and serving phases, which is a common issue in machine learning engineering systems (Zinkevich, 2010).
@@
Many ranking systems in the literature rely on pre-trained items and customer embeddings. Our experiments reveal that similarly to NLP tasks (Devlin et al., 2019), the effectiveness of our models significantly increases when the pre-trained input item embeddings are further fine-tuned on the ranking task. Notably, we show that if these embeddings are not continuously trained, the candidate generation model shows substantially less customer engagement. To address the item cold-start problem, we introduce epsilon-greedy exploration by blending fresh items from additional candidate sources into the organic ranking. To address the customer cold-start problem, we leverage customer context and in-session data.
@@ 2. RELATED WORK
Thanks to their advantages over traditional deep-learning-based models, sequence-based recommender systems (Kang and McAuley, 2018; Sun et al., 2019) in their two flavors of language modeling (CLM and MLM) (de Souza Pereira Moreira et al., 2021; Petrov and Macdonald, 2023) have gained wide traction. These systems have proven powerful in modeling customer behavior as a sequence of actions due to their capability to 1) capture both short-term and long-term interests (Xia et al., 2023); and 2) the ability to compute complex feature interactions. However, most existing works are trained and evaluated on public datasets that sometimes are not adequate for sequential recommendation tasks (Hidasi and Czapp, 2023; Kang and McAuley, 2018; Sun et al., 2019; de Souza Pereira Moreira et al., 2021). Moreovewer, most of the works focus only on offline experiments, with only a few works reporting actual customer impact through end-to-end A/B testing in large-scale environments (Xia et al., 2023; Pancha et al., 2022). Our paper extends this line of work and demonstrates the usefulness of these models in real-world applications that include personalized item browsing and search. Unlike related approaches (Chen et al., 2019; Xia et al., 2023), our approach to scoring candidate items does not suffer from performance issues caused by a long sequence length due to concatenating the candidate item embeddings with the customer action sequence embeddings as in (Chen et al., 2019) or limiting the input embedding size as in (Xia et al., 2023). We do not observe a significant drop in ranking performance when feeding only the average embedding of the candidates as a fixed input position into the transformer network compared to including all candidate items as part of the input sequence.
Moreover, only a few published studies (Steck et al., 2021; Celikik et al., 2022a; Chen et al., 2019; Xia et al., 2023; Celikik et al., 2022b) include heterogeneous inputs such as context and content-based features, which help address the cold-start problem and address data sparsity by improving generalization. For example, (Chen et al., 2019) considers customer profiles and contextual features by using wide & deep learning (WDL), which relies on a concatenation of signals in the output of the network, making it inadequate to capture powerful feature interactions. (Xia et al., 2023) employs deep and cross-network (DCN) on top of a transformer network to explicitly model feature-crosses, which significantly increases the number of parameters of the model.
@@ 3. SYSTEM DESIGN
Designing a highly performant, scalable, and steerable ranking platform entails challenges and complex choices. The existing literature often focuses only on subsets of them (Lakshmanan et al., 2020; Covington et al., 2016; Yang et al., 2020; Essinger et al., 2021), while this section aims to navigate through them holistically. We describe foundational design principles and provide an overview of our system architecture and components (see Figure 2). The presented design is generalizable and applicable to other retrieval and ranking use cases and setups.
@@
Composability and orchestration of multiple models. Our platform enables the combination of various models for several applications. This can be done ”vertically”, by layering them on top of one another for multi-layered ranking and retrieval which allows scalability and avoids separation of concerns. Ranking models can also be combined ”horizontally”, by blending the outputs of multiple models or candidate generators. Our platform consists of three layers. The first layer retrieves relevant candidates from multiple candidate generators, possibly generating different content types (e.g. items, outfits, stories, etc.) for use cases such as feeds. Each candidate generator typically entails a (lightweight) ranking model. Subsequent layers refine these selections. The ranking layer applies heavy personalized models for ranking pre-selected items of possibly different types. The policy layer ensures compliance with business or product specifications. Mixing strategies mix outputs to suit specific needs, such as combining different content types in desired proportions or balancing popular, fresh, and personalized content (Gomez-Uribe and Hunt, 2016). Model blending is also the blueprint for building ranking ensembles, where outputs of multiple models are combined either by score weighting or meta rankers.
@@
Scalable platform. As already mentioned, a multi-layered model architecture allows high scalability. The more accurate, but computationally heavier part of the ranking is performed on later layers only on a small subset of candidates obtained from the less accurate but more computationally efficient candidate generator layers. Thus, the highest scalability requirements are placed on the candidate generator. Scalability of the candidate generator layer is achieved through a vector store allowing efficient indexing, scoring, and retrieval capabilities to scale for a large and growing item catalog and customer base. Our platform includes infrastructure to compute customer embeddings in near real-time each time a customer accesses the platform. Item embeddings are computed in a streaming-based fashion whenever a new item is introduced and afterward asynchronously indexed.
Shared near real-time serving infrastructure. Thanks to the similarity of model types and architectures employed in the candidate generation and ranking layers, the training dataset and a large part of our serving infrastructure are reused across the layers. In addition, the use of deep sequence models allows for lightweight feature engineering pipelines consisting of embedding mappings. These mappings reside in the model graph, which guarantees efficiency and consistency between training and serving. Our online feature store is shared between the different ranking layers allowing effective caching of inputs for a given ranking request as well as decreasing the engineering complexity.
@@
Steerable ranking. Our system’s flexibility allows for external adjustments to ranking objectives via multi-objective optimization to align with business goals such as customer engagement and revenue. It supports ranking and mixing diverse content types through its candidate generators and mixing components. Finally, business heuristics are applied in the policy layer.
4. CANDIDATE GENERATION LAYER
@@
The objective of our candidate generation layer is to generate personalized item candidates from the item catalog for each individual customer efficiently in near real-time by scoring a vocabulary of millions of items. According to our findings, a personalized and context-aware candidate generator is essential for the performance of the overall ranking system. The top-500 candidates are then reranked by a heavy ranking model in the ranking layer.
@@
We follow the classical two-tower approach (Yi et al., 2019), where the customer tower processes historical customer action sequences and contextual data to generate a customer embedding while the item tower is responsible for generating item embeddings. These embeddings are then combined by using dot product to generate a score per item as shown in Figure 3 (a theoretical justification about the expressiveness of the two-tower model is provided in the Appendix, subsection 10.3).
@@
To compute a customer embedding from the customer action sequence and the context, we employ a transformer encoder and use causal language modeling (CLM) as in (Kang and McAuley, 2018), processing each customer sequence once per epoch. We predict the subsequent item in the sequence while preventing backward attention by using a causal mask. Scores are computed by multiplying the output vector of the encoder with an output item embedding matrix.
Although trained together, the item and customer towers are deployed and operated independently. The item tower generates item embeddings that are indexed in a vector store. The streaming-based indexing takes place whenever a new item joins the platform, the model is retrained or the inputs of an existing item have been changed. The customer tower is invoked to generate customer embeddings each time a customer accesses the platform. This follows a call to a feature store to fetch all customer-specific inputs, such as context and behavioral sequence. These are then stored in a cache for other transformer-based models in the funnel to reuse. The freshly computed customer embeddings are then used to find items with similar embeddings in the nearest-neighbors index.
For simplicity, items in the item tower in Figure 3 are represented by using a single embedding that jointly encodes product metadata (brand, category, material, etc.) and visual cues. It is worth noting that in this common setting, trainable input embeddings on the ranking task in the candidate generation layer performed substantially better in both offline and online tests. Details on how different types of input signals from the customer journey data (e.g. contextual and customer action sequence data) are encoded in the customer tower of the model are captured in subsection 5.2.
@@ 5. RANKING LAYER
The objective of the ranking layer is to rank items returned by the candidate generation phase by their relevance to the customer and their context by using a powerful ranking model. We model this task as a pointwise multi-task prediction problem, where we predict the probability of the customer performing any of the following positive actions on a candidate item: click, add-to-wishlist, add-to-cart, purchase given the context and their past behavioral data. If a candidate item is associated with any of these positive actions, we consider it a positive item, otherwise, it is considered a negative.
@@ 5.1. Model Architecture
Figure 4 depicts the architecture of the model in the ranking layer. It consists of four main parts: embedding layer, item candidate embedding, customer-context embedding computed via a self-attention mechanism, a prediction head for each of the target action types, and a shallow position branch used for position debiasing.
@@
For each of the target actions, we define a prediction head, which takes customer and candidate item representations as inputs. Similar as in the candidate generation model, the score per target action is obtained by computing a dot product between the customer-context embedding and all candidate item embeddings in parallel, after passing them through a FFN. A sigmoid function is used to normalize the score and interpret it as a probability. During training, each prediction head contributes equally to the loss, while at serving we produce the final ranking by weighting the scores of each prediction head. The weights are dynamically configurable and determined analytically depending on the customer touch point.
@@
The training objective of our main candidate generator is based on predicting the next customer action, however, our data training pipelines are configurable to predict actions in a longer future time window by masking actions in immediate distance to balance long and short-term customer preferences and improve diversity. A simple masking heuristic we employed from (Xia et al., 2023) called ”random time window mask” proved effective and significantly increased the number of explored item categories in an A/B test.
@@ 5.2. Encoding of the Customer Journey Data
In this section, we elaborate on how we encode holistic customer journey data into models in both layers. This consists of (1) behavioral data which includes customer action sequences, item and action metadata, and temporal data, and global and local contextual data. We argue and demonstrate in our experiments that providing complete and heterogeneous information to the model is important for predictions that are personalized and contextually relevant.
Behavioural data. As already mentioned, customer action sequences are encoded using transformer encoders as these have been proven more effective than other approaches (Xia et al., 2023). Each action in an action sequence is represented by item embedding that encodes domain-specific visual information, categorical item metadata, action type, and timestamp embeddings. The categorical item metadata consists of relevant attributes such as brand, color, pattern, category, material, etc. The timestamp embedding encodes quantized timestamps as measured from the beginning of the model training. Temporal data is crucial for modeling customer behavior across sessions. Since customer intent can vary drastically across different sessions, modeling action sequences while ignoring this structure affects performance negatively (Rahmani et al., 2023; Celikik et al., 2022b). All inputs are passed through trainable embedding to project discrete or bucketized values into low-dimensional spaces. Item and action-specific embeddings are concatenated with the item embedding.
@@ 5.3. Position Debiasing
In the context of ranking systems, feedback loops occur when a model influences customer interactions which can lead to biased relevance data. Typically, items ranked higher by the model receive more customer attention, causing position bias. Position bias causes a skewed representation of actual customer preferences which in turn may amplify and degrade model performance over time due to the feedback loop. To address this, we incorporate position information as a feature into a model to separate the effect of the position and the true relevance of the probability with which the customer would interact with an item. A debiased model is conditioned on positions during training (right branch in section 5) and position-independent during serving by setting positional feature to fixed values to counter position bias (Zhao et al., 2019). The position branch is separate from the rest of the model due to the asymmetry between training and serving. We performed additional A/B test which confirmed that adding position debiasing resulted in increasing long-tail utilisation by 5.7% (measured by item popularity at top-6) as well as improved catalog item utilization by 3.1% (measured by effective catalog size at top-10) while not deteriorating engagement and financial metrics.
@@ 6. POLICY LAYER
The last stage of the system, the policy layer, is responsible for the final page composition. Here, multiple re-ranked candidate items are combined into one, performing granular, page-level optimization, and applying heuristics, business rules, and filters depending on the use case. In the following, we describe how we promote fresh (cold-start) items while simultaneously introducing exploration into the system. We also describe some common heuristics applied in this layer to meet product requirements.